### v2

# Enforcing MAVLink Safety & Security Properties Via Refined Multiparty Session Types

Abstract
A compromised system component can issue message sequences that are legal while also leading the overall system into unsafe states. Such stealthy attacks are challenging to characterize, because message interfaces in standard languages specify each individual message separately but do not specify safe sequences of messages. We present initial results from ongoing work applying refined multiparty session types as a mechanism for expressing and enforcing proper message usage to exclude unsafe sequences. We illustrate our approach by using refined multiparty session types to mitigate safety and security issues in the MAVLink protocol commonly used in UAVs.
Keywords:
Integrated formal methods Dynamic checking DATUM1 Introduction
Attackers who are already inside a system can mount stealthy attacks in which they send apparently innocuous messages that may still cause damage. We aim to limit their damage by dynamically checking that messages conform to a protocol, enabling critical systems to have verified safety guarantees.
One approach to keeping attackers out of critical systems is exemplified by the DARPA High-Assurance Cyber Military Systems (HACMS) program [11]. HACMS combines SMACCMPilot [13] and seL4 [18] to prevent external adversaries from compromising the autopilot system of a UAV.
However, researchers have since identified novel attacks against UAVs [17, 16]. The main such attack is a stealthy attack, in which a compromised ground control station (GCS) sends seemingly innocuous messages to the UAV. The GCS might, for example, be following the dictates of a disgruntled employee. Such messages take advantage of an incorrectly designed or implemented protocol to evade intrusion detection and still cause unsafe behavior. As a response to stealthy attacks, researchers have proposed runtime verification [24, 25, 2].
Our approach, which we call Dynamically Assured Typed Universal Messaging (DATUM), checks that the communication between the controller and the controlled system (e.g., the GCS and the UAV) conforms to a specified protocol. The protocol itself can also be checked to validate that it preserves safety.111We do not discuss the details of checking the safety of protocols in this paper.
In this paper, we formalize a substantial portion of the MAVLink protocol [20] used by GCSs to control UAVs. MAVLink specifies more than 200 messages and 100 custom data types. From the MAVLink specification, we generated more than 8k lines of code.
Our approach benefits two audiences. First, operators can use our approach to mitigate the risks posed by stealthy attacks while they await patches. Second, engineers can validate a patch’s modifications before releasing it.
To summarize our contributions, we:
Describe a methodology that can prevent stealthy attacks against UAVs.
Formalize a substantial part of the MAVLink protocol [20] using DATUM.
2 Background
2.1 Session Types
DATUM uses session types, a rich typing discipline [14]. Binary session types can enforce constraints, such as message ordering, and capture temporal properties between two actors, such as the absence of deadlocks [8]. Multiparty session types (MPSTs) extend these properties to systems with multiple actors [29, 15]. For each exchanged message, MPSTs can express the message sender, receiver, and payload type. MPSTs have led to many real-world applications [22, 5, 30].
However, MPSTs cannot reason about payload values, as they cannot constrain the values being exchanged in the systems. For example, MPSTs cannot write a protocol where sends a number , and then sends that same to . Refined multiparty session types (RMPSTs) [28, 33, 32] can specify such constraints by expressing properties of values within messages. While RMPSTs are often used in static reasoning; we are the first to use them for dynamic checking.
An RMPST has a syntax that follows one of the cases given below:
Messages: A message specifies a sender (e.g., ), a receiver (e.g., ), a label (e.g., MSG), and the message’s payload (e.g., ). The payload is specified with a type (e.g., ) and may optionally have a refinement, (e.g., ).
Guarded Choice: A refinement may contain a choice that depends on values. The following is an example where sends to different messages depending on the refinement:
Mu-Recursion: A recursive protocol can be specified using the operator. For example, the following recursive protocol has as its recursion condition, where is incremented after every recursion.
End: The End construct specifies a protocol’s endpoint. The overall protocol terminates when all participants reach their endpoints. For example:
2.2 The F* Interactive Theorem Prover
F* is an interactive theorem prover that integrates the Z3[9] SMT solver for automated proofs, while also allowing users to write constructive proofs[26]. This combination allows trivial properties to be proved with minimal input, while allowing for interactive proofs of other properties. Both kinds of properties can be proven to hold in all possible executions.
We chose to represent our RMPSTs in F* for its proof automation and native support for refinements. Previous work showed promise in the formalization of a binary-session-typed calculus in theorem provers [31, 21]. Other works showed how to prove properties about session subtyping [10]. Zooid [6] is a domain specific language embedded in Coq [7] that allows users to define protocols using MPSTs. Zhou et. al. showed how an F* API can be used to statically verify an RMPST protocol specification [33]. DATUM allows for protocols to be described as F* terms, allowing runtime traces to be checked.
3 Methodology
3.1 Case Study I: Resource Usage
We demonstrate our methodology for preventing stealthy attacks against the MAVLink mission sub-protocol. MAVLink is the de facto standard for GCS-UAV communications despite numerous safety and security shortcomings [1, 12, 19, 27].
MAVLink allows the GCS to send mission directives to the UAV. Mission directives include waypoints, geofences, and safepoints. The GCS starts the mission sub-protocol by sending the MISSION_COUNT MAVLink message, where is the number of mission items the GCS will send. Consequently, the UAV expects exactly mission items.
Issue reports show how a UAV running ArduPilot accepted fewer mission items than its declared mission count [3]. This error was caused by a failed waypoint upload from an earlier flight. These uploads were kept in a buffer until the next use of the mission sub-protocol. Midway through the mission, the UAV began following a previous mission’s flight plan. Such unexpected changes in the flight plan could have compromised the safety of the UAV. The mission sub-protocol should have the capability to reject missions based on missing items.
3.2 Mitigation With RMPSTs
In our methodology, the sub-protocol initiated by the message MISSION_COUNT is specified as an RMPST (shown below) with recursion over guarded by the refinement . The recursive protocol offers a choice between the messages MISSION_REQUEST_INT and MISSION_ACK. The MISSION_REQUEST_INT message must have a payload between and and the GCS must request missions sequentially. We enforce the second condition by comparing the payload to , which increments after every iteration. Following the MISSION_REQUEST_INT message, the UAV sends a MISSION_ITEM_INT message and recurs while incrementing the variable.
The sub-protocol ends with a MISSION_ACK message and a MISSION_RESULT payload . When an error occurs, the UAV sends a MISSION_ACK message with the respective ERROR payload. If all missions upload successfully, reaches value and the MISSION_ACK message represents a correct sub-protocol run.
3.3 Using DATUM To Write RMPSTs in F*
Figure 1 shows the RMPST from our case study formalized in F*. This listing relies on the DATUM framework that we have developed. The helper type mission_req shows how we can represent the refinement on the MISSION_REQUEST_INT GCS message that appears in the case study. Other helper types like mission_item are similar. Figure 1 excludes their definitions for brevity.
Notably, DATUM does not utilize F*’s native refinement types when writing an RMPST. Utilizing F*’s native refinement types would require expensive dynamic type-checking of the entire F* programming language to attest that a communication conforms to an RMPST. Instead, DATUM provides the RefType constructor, shown in Figure 1. This reduces the dynamic attestation problem to checking that a predicate evaluates to true.
DATUM provides sound support for modeling RMPSTs. In DATUM’s model, RMPSTs have four constructors: Offer, Mu, Recur, and End. Both Offer and Mu introduce new variables. New variables are represented as function parameters.
DATUM uses F*’s effect system to prevent function parameters representing RMPST variables from creating ill-formed RMPSTs. F*’s effect system is an extensible lattice of monadic effects [26]. DATUM defines a hidden monad that prevents variables from being used outside of refinements. A value of type hidden cannot be used in F*’s default effect Tot. However, DATUM can use it when performing I/O during dynamic attestation.
DATUM automatically generates a dynamic attestation tool from the specification of MAVLink, as shown in Figure 2. When using DATUM, communication between the GCS and the autopilot software passes through a MAVLink proxy. The MAVLink proxy uses the attestation tool to check each message. Non-compliant messages identified by DATUM are reported to the UAV’s pilot.
3.4 Inside of DATUM
Figure 2 shows an overview of our approach. First, we automatically generate F* types from MAVLink’s machine-readable specification file. Then, we use DATUM to write an RMPST describing MAVLink. Inside the MAVLink RMPST, we include refinements that mitigate vulnerabilities.
The MAVLink specification [20] defines more than 300 types to describe messages and their payloads. These types must be written in F* to use DATUM to define RMPSTs. Manually writing these 300 types in F* would be tedious. Instead, we use automation to generate F* types. Consider the MISSION_COUNT message, whose F* and DATUM representation is shown in the left call-out of Figure 2. This code is automatically generated from the common.xml file that MAVLink distributes. The right call-out in Figure 2 shows the part of common.xml that contributes to the generation of the mission_count type. In total, automation generates lines of F* code.
3.5 Additional Case Studies
In this section we discuss two additional case studies and demonstrate their respective mitigation using RMPSTs. For brevity, we omit their formalization.
3.5.1 Case Study II: Overly Permissive Protocol
An overly permissive protocol bug occurs when a protocol allows values that may exceed safe ranges. The PX4 autopilot software includes over 200 configurable parameters. These parameters are used by PX4 to navigate the UAV. Some parameters have interdependencies; that is, the safe range of a parameter may depend on the values of related parameters. For example, in PX4 the MC_PITCHRATE_FF and MC_PITCH_P parameters represent the feed-forwarding pitch rate and pitch gain. When both parameters are set to high values, the UAV makes jerky movements to try to stabilize itself, but such movements can result in loss of stability. In these situations, a low maximum pitch rate (MC_PITCHRATE_MAX) value is needed to ensure stability.
Overly permissive protocol bugs can lead to scenarios where the UAV crashes due to communications that do not violate the parameter’s range requirements. This vulnerability could be used by a stealthy attacker. Such an attacker could send malicious messages to the UAV to change these parameters with the message PARAM_SET, while still following the MAVLink protocol [17]. However, RMPSTs can constrain dynamic parameter changes with respect to value dependencies. For example, this value dependency can be expressed as follows (where and are weights for the parameters involved):
| GCS UAV: | |||
The dynamic attestation checker applies the value dependency predicate in the refinement to the message payload . Thus DATUM enforces that the PARAM_SET message is only seen by the UAV if it is safe to execute.
3.5.2 Case Study III: Precondition Violations
Some MAVLink messages, when executed at the wrong time, can cause harm to the UAV, however message preconditions are not enforced. One such message, MAV_CMD_DO_PARACHUTE, deploys the parachute in ArduPilot. The documentation defines the necessary preconditions for this message: the UAV is not ascending, the payload equals 2, its motors are armed, it is not in ACRO or FLIP mode, and its altitude is above CHUTE_ALT_MIN parameter. Deploying the parachute while violating any of the preconditions, even though they are not enforced, can lead to a crash. Thus, an attacker could wait for an opportunistic moment to deploy the parachute [16].
DATUM can enforce these precondtions using RMPSTs, as shown below.
4 Evaluation
We evaluated our methodology by retrofitting systems equipped with ArduPilot 4.5.7 [4] and PX4 1.15.2 [23]. ArduPilot and PX4 were evaluated using their software in the loop (SITL) systems. All experiments were conducted on a machine equipped with an Intel i9 processor, 64 GB of RAM, and Ubuntu 24.04.
In our evaluation, we simulated a mission with 100 waypoints, each uploaded by the GCS to the autopilot software. Our experiments measured memory usage and message latency.222 To measure message latency, we recorded the timestamps when each message was sent and received. This required modifying the flight control software to record the appropriate timestamps. The results of our experiments are shown in Table 1.
| Software | RSS | RSS w/ Datum | Latency | Latency w/ Datum |
| ArduPilot | 8,096 KB | 21,824 KB | 29.60 6.09 | 56.74 6.75 |
| PX4 | 18,490 KB | 32,208 KB | 20.96 1.97 | 57.70 6.53 |
The overhead numbers are small when read as absolute figures. However, relative to the values without dynamic attestation the overhead seems high. We note that the relative overhead would be smaller in an actual deployment where the communication latency would be much greater than in our evaluation. (Our evaluation setup consisted of a single machine with communications exchanged on its loopback interface. Thus, communication latency was small, less than ; however, latency between the GCS and UAV is typically 2-5 ms.)
5 Conclusion
In this work, we have shown how to use lightweight formal methods to retrofit existing systems to mitigate real-world safety and security weaknesses in a widely deployed protocol. Our methodology, DATUM, uses RMPSTs to write protocols with logical constraints. These constraints are dynamically checked. DATUM allows GCS software to prevent attacks that use unpatched bugs and to fine-tune protocols for future patches.